OpenStudy (anonymous):

i am having trouble with this : Let f(x) = x – 2 and g(x) = 3x + 4. What is f(g(x)) ?

OpenStudy (saifoo.khan):

f(x) = x-2 f(g(x) = 3x+4 -2 DYS.

OpenStudy (anonymous):

plug in g(x) for every x in f(x) f(g(x)) = (3x+4)-2 f(g(x))=3x+2

OpenStudy (anonymous):

thats exactly what i got i just wasnt 100% sure